OpenStudy (anonymous):

The function f(x)=4(2)^x represents the growth of a butterfly population every year in a remote swamp. Jan wants to manipulate the formula to an equivalent form that calculates five times a year, not just once a year. Which function is correct for Jan's purpose, and what is the new growth rate?

OpenStudy (anonymous):

A. f(x)=4(1.15)^x ; growth rate is 5% B. f(x)=4(1.15)^5x ; growth rate is 115% C. f(x)=4(2)^x ; growth rate is 200% D. f(x)=4(2)^x ; growth rate is 5%

zepdrix (zepdrix):

Hey :)

zepdrix (zepdrix):

They want an equation that is `equivalent` to the first one given. So we'll have to start with that, and manipulate it to get a 5 in the exponent.

zepdrix (zepdrix):

I want to change the 2 in a fancy way, applying some exponent rules,\[\Large\rm \color{orangered}{2}=2^1=2^{(\frac{1}{5}\cdot5)}=\color{orangered}{\left(2^{1/5}\right)^5}\]

zepdrix (zepdrix):

So our equation\[\Large\rm y=4\left[\color{orangered}{2}\right]^x\]can be written as\[\Large\rm y=4\left[\color{orangered}{\left(2^{1/5}\right)^5}\right]^x\]

zepdrix (zepdrix):

Applying exponent rules a little further\[\Large\rm y=4\left(2^{1/5}\right)^{5x}\]gives us the 5 in the exponent like we want. We need to recalculate the base of this exponential though.

zepdrix (zepdrix):

2^(1/5) is probably 1.15 since that's what is showing up in your answer choices. Then you need to think about what growth rate 1.15 corresponds to. 5% or 115%?

OpenStudy (anonymous):

115?

OpenStudy (anonymous):

@zepdrix are you still there?

zepdrix (zepdrix):

115% ? Yah that sounds right! :)

OpenStudy (anonymous):

I already had B selected, I was just double checking. Word problems aren't really my thing. Thank you! I will fan and give you a medal :)

zepdrix (zepdrix):

yay team \c:/